OpenStudy (anonymous):

How many different arrangements of the letters in the word "PURPLE" are there? a)30 b)360 c)720

jimthompson5910 (jim_thompson5910):

how many letters are there

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

so if you had 6 distinct letters, then you would have 6! = 6*5*4*3*2*1 = 720 different ways to arrange them BUT you have two copies of P that are indistinguishable, so we double counted when we got 720, which is why we must divide by 2 to get 720/2 = 360

OpenStudy (anonymous):

I see thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw